By what potential difference must an electron (m = 9.11*10^-31 kg) be accelerated to have a wavelength of 5.16*10^-12 m? i keep getting wied answers, thanks in advance.
Answer #1

For an electron accelerated by potential difference V the de broglie wavelength is related to V as

\lambda = \frac{h}{\sqrt{2meV}} .......... 1

we have

h = 6.6 *10^{-34} Js

m = 9.11 * 10^{-31} kg

e = 1.6 * 10^{-19} C

\lambda= 5.16 * 10^{-12} m

Substituting these values in equation 1

5.16 * 10^{-12}= \frac{1.227 *10^{-9}}{\sqrt{V}}

{\sqrt{V}} = 237.8

V = 5.65*10^4 V

We are given